# Data-driven decision-making under uncertainty with entropic risk measure

### v1

@@ A.3 Proof of Theorem 2
$\ell_{*}({{\boldsymbol{z}}},{{\boldsymbol{\varphi}}})=\left\{\begin{array}[]{cl}0&\text{if }{{\boldsymbol{\varphi}}}={{\boldsymbol{z}}}\\ -\infty&\text{otherwise}.\end{array}\right.$ (30)
Thus, problem (27) reduces to:
@@
subject ${\rm subject\mbox{ }\rm to}\quad$ $\hat{{{\boldsymbol{\xi}}}}_{i}^{\top}{{\boldsymbol{z}}}+\epsilon\|{{\boldsymbol{z}}}\|_{*}\leq t_{i}\,,\;\forall i\in[N].$ (32)
Replacing the epigraph variable ${{\boldsymbol{t}}}$, we obtain that the objective reduces as follows:
$\frac{1}{\alpha}\log\left(\frac{1}{N}\sum_{i=1}^{N}\exp(\alpha(\hat{{{\boldsymbol{\xi}}}}_{i}^{\top}{{\boldsymbol{z}}}+\epsilon\|{{\boldsymbol{z}}}\|_{*}))\right)$ $\ =\ \frac{1}{\alpha}\log\left(\exp(\alpha\epsilon\|{{\boldsymbol{z}}}\|_{*}))\cdot\frac{1}{N}\sum_{i=1}^{N}\exp(\alpha\hat{{{\boldsymbol{\xi}}}}_{i}^{\top}{{\boldsymbol{z}}})\right)$

### v3

@@ A.3 Proof of Proposition 3
[…]
Moreover, ${\mathbb{E}}[L({{\boldsymbol{\xi}}})^{2(1+1/\varsigma)}]$ is finite based on Assumption 3(A.5), which allows us to conclude that ${\mathbb{E}}[\kappa({{\boldsymbol{\xi}}})^{2}]<\infty$.
We now have all the components to apply Jiang et al., (2020, Theorem 3.2) which states that for any $\gamma>0$, there exists an $A>0$, independent of $N$, such that:
${\mathbb{P}}\left(\mathop{\rm sup}_{{{\boldsymbol{z}}}\in\mathcal{Z}}\left|{\mathbb{E}}_{\hat{{\mathbb{P}}}_{N}}[\exp(\alpha\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}}))]-{\mathbb{E}}_{{\mathbb{P}}}[\exp(\alpha\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}}))]\right|\geq A/\sqrt{N\gamma}\right)\leq\gamma,$ (22)
for sufficiently large $N$, thus concluding that the empirical exponential utility converges uniformly to the true exponential utility, with a rate of convergence of $\mathcal{O}(1/\sqrt{N})$.
Next, we show that empirical risk converges uniformly to the true risk by exploiting the properties of the $\log$ function, namely that $\log(1+\epsilon)\leq\epsilon$ for all $\epsilon\geq 0$, and $\log(1-\epsilon)\geq-\epsilon/(1-1/e))$ for all $\epsilon\in[0,1-1/e]$. This result is summarized in Lemma 12 and can be pictorially verified in Figure 13.
(a) $\log(1+\epsilon)\leq\epsilon,\;\forall\epsilon\geq 0$.
(b) $\log(1-\epsilon)\geq-\epsilon/(1-1/e)\;\forall\epsilon\in[0,1-1/e]$
Figure 13: Plot of the inequalities that the $\log$ function satisfies around 00.
In the subsequent steps of the proof, we consider $N$ to be large enough such that $\frac{A}{\sqrt{N\gamma}\exp(\alpha\rho^{*})}\leq 1-1/e$, with $\rho^{*}:=\mathop{\rm min}_{{{\boldsymbol{z}}}\in\mathcal{Z}}(1/\alpha)\log({\mathbb{E}}_{{\mathbb{P}}}[\exp(\alpha\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}}))])$ and focus on scenarios for which the absolute difference between the empirical utility and true utility in (22) is upper bounded by $A/(\sqrt{N\gamma})$ for all ${{\boldsymbol{z}}}\in\mathcal{Z}$. Specifically, we will show that it implies that the absolute difference between the empirical risk and true risk is upper bounded by $A/(\alpha\sqrt{N\gamma}\exp(\alpha\rho^{*}))$ for all ${{\boldsymbol{z}}}\in\mathcal{Z}$. To simplify notation, we further denote by $\hat{u}_{N}({{\boldsymbol{z}}}):={\mathbb{E}}_{\hat{{\mathbb{P}}}_{N}}(\exp(\alpha\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}})))$, $u({{\boldsymbol{z}}}):={\mathbb{E}}_{{\mathbb{P}}}[\exp(\alpha\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}}))]$, $u^{*}:=\mathop{\rm min}_{{{\boldsymbol{z}}}\in\mathcal{Z}}{\mathbb{E}}_{{\mathbb{P}}}(\exp(\alpha\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}})))=\exp(\alpha\rho^{*})$. For any $\gamma>0$, from (22), we have:
$\begin{array}[]{lll}&\lvert\hat{u}_{N}({{\boldsymbol{z}}})-u({{\boldsymbol{z}}})\rvert<\frac{A}{\sqrt{N\gamma}}&\forall{{\boldsymbol{z}}}\in\mathcal{Z}\\ \implies&\frac{\lvert\hat{u}_{N}({{\boldsymbol{z}}})-u({{\boldsymbol{z}}})\rvert}{u^{*}}<\frac{A}{\sqrt{N\gamma}\exp(\alpha\rho^{*})}&\forall{{\boldsymbol{z}}}\in\mathcal{Z}\\ \implies&\frac{\lvert\hat{u}_{N}({{\boldsymbol{z}}})-u({{\boldsymbol{z}}})\rvert}{u({{\boldsymbol{z}}})}<\frac{A}{\sqrt{N\gamma}\exp(\alpha\rho^{*})}&\forall{{\boldsymbol{z}}}\in\mathcal{Z}\\ \implies&\frac{\hat{u}_{N}({{\boldsymbol{z}}})}{u({{\boldsymbol{z}}})}<1+\frac{A}{\sqrt{N\gamma}\exp(\alpha\rho^{*})}&\forall{{\boldsymbol{z}}}\in\mathcal{Z}\\ \implies&\frac{1}{\alpha}\log\left(\frac{\hat{u}_{N}({{\boldsymbol{z}}})}{u({{\boldsymbol{z}}})}\right)<\frac{1}{\alpha}\log\left(1+\frac{A}{\sqrt{N\gamma}\exp(\alpha\rho^{*})}\right)&\forall{{\boldsymbol{z}}}\in\mathcal{Z},\end{array}$
where in the first implication we divided both sides by $u^{*}$, while the second implication follows since $u^{*}\leq\mathop{\rm min}_{{{\boldsymbol{z}}}\in\mathcal{Z}}u({{\boldsymbol{z}}})$. The last implication follows by applying the logarithm on both sides of the inequality and dividing by $\alpha$. The last expression can be further simplified since by the properties of the logarithm function, we have $\log(1+\epsilon)\leq\epsilon$ for all $\epsilon\geq 0$, hence: